OpenStudy (anonymous):

Find the three geometric means between 15 and 1215

OpenStudy (anonymous):

@nikato

OpenStudy (nikato):

sorry about that. do u still need help? @StaceyMarie

OpenStudy (nikato):

well whether u finish ur problem or not. u have 15,_,_,_,1215 15 is the 1st term and 1215 is the 5th plug it into the geometric sequence formula n=ar^n-1 where n=term a=first term of sequence r=ratio n=# of terms to solve the problem the first thing to do is to find r

OpenStudy (nikato):

so plug in 15 and 1215 and divide the 2 to find r 1215=15r^4 ------------- 15 =15r^0 so 81=r^4 r=3 so to find the term term just multiply the previous term by 3. do this 3 times to fidn the three geometric means

OpenStudy (anonymous):

Oh thank you but I figured it out. it was actually pretty simple when I understood it a bit better

OpenStudy (nikato):

no problem. and yea, thsi problem wasnt that difficult, just lengthy